A solution of 77% vinegar is mixed with a solution of a 25% vinegar to form 312 liters of a 64% solution . How many liters of the 77% solution must be used ?
Answer by stanbon(75887)

A solution of 77% vinegar is mixed with a solution of a 25% vinegar to form 312 liters of a 64% solution . How many liters of the 77% solution must be used ?
------
Equation:
0.77x + 0.25(312-x) = 0.64*312
----
77x + 25*312 - 25x = 64*312
----------
52x = 39*312
x = 234 liters (amt. of 77% solution needed)
----
312-234 = 78 liters (amt. 25% solution needed)
===================
Cheers,
Stan H.
===================